### v2

# How Well Do LLMs Handle Cantonese? Benchmarking Cantonese Capabilities of Large Language Models

Abstract
The rapid evolution of large language models (LLMs) has transformed the competitive landscape in natural language processing (NLP), particularly for English and other data-rich languages. However, underrepresented languages like Cantonese, spoken by over 85 million people, face significant development gaps, which is particularly concerning given the economic significance of the Guangdong-Hong Kong-Macau Greater Bay Area, and in substantial Cantonese-speaking populations in places like Singapore and North America. Despite its wide use, Cantonese has scant representation in NLP research, especially compared to other languages from similarly developed regions. To bridge these gaps, we outline current Cantonese NLP methods and introduce new benchmarks designed to evaluate LLM performance in factual generation, mathematical logic, complex reasoning, and general knowledge in Cantonese, which aim to advance open-source Cantonese LLM technology. We also propose future research directions and recommended models to enhance Cantonese LLM development [^1] The code and data are available anonymously github. The anonymously github link is as follow: https://github.com/jiangjyjy/Yue-Benchmark.
How Well Do LLMs Handle Cantonese? Benchmarking Cantonese Capabilities of Large Language Models
Jiyue Jiang, Pengan Chen, Liheng Chen, Sheng Wang, Qinghang Bao, Lingpeng Kong, Yu Li, Chuan Wu The Chinese University of Hong Kong, The University of Hong Kong jiangjy@link.cuhk.edu.hk {cpa2001, clh648, u3009618, bill6176}@connect.hku.hk, lpk@cs.hku.hk liyu@cse.cuhk.edu.hk cwu@cs.hku.hk
Figure 1: This is number of publications in the ACL Anthology indexed by languages as of September 2024. Following Xiang et al. (2024), we retrieve the publications via searching the language name in either the title or the abstract from the ACL Anthology.
1 Introduction
Increasingly impactful and LLMs have emerged (e.g., GPT-X, Llama-X, etc.), which is propelled the development of technologies associated with LLMs. As shown in Figure 1, NLP research has predominantly concentrated on creating models for English and a few other languages that have substantial data resources Aji et al. (2022). The scarcity of data is often identified as the primary obstacle impeding advancements in NLP for languages that are less represented Hu et al. (2020); Joshi et al. (2020); Aji et al. (2022), particularly for LLM-related technologies.
Figure 2: Overview of the paper: We begin by summarizing approaches from small-scale neural networks in Cantonese, then progress to LLMs (work involving existing Cantonese LLMs). In these LLMs, researchers place a greater emphasis on alignment compared to pre-training. Consequently, we introduce four new benchmarks and a translation datatset to evaluate the Cantonese capabilities of LLMs. We analyze the performance of mainstream LLMs on these benchmarks and, in combination with the inherent challenges of Cantonese itself, identify three insightful research opportunities, and we summarize the models that perform good for each specific task. (Figure 5).
Cantonese (Yue language), spoken by over 85 million people worldwide Xiang et al. (2024), has seen slower technological development, particularly in the LLMs. Language technologies for Cantonese have not yet reaped the benefits of this revolution Xiang et al. (2022). As indicated in Figure 1 and Table 1, there is a low number of recent research publications related to Cantonese, especially when compared to the population ratio. Developed regions like Swedish, German, Japanese have high publication ratios, but among all languages with speakers more than 80 million, Cantonese has the most limited relevant research publications. Given that the Guangdong-Hong Kong-Macau Greater Bay Area is one of the most economically vibrant regions in the world [^2] https://www.bayarea.gov.hk/filemanager/en/share/pdf/Outline_Development_Plan.pdf and that many countries (e.g., Singapore, Malaysia, Australia, Canada, U.S., etc.) have a large Cantonese-speaking population, advancing Cantonese LLM technology represents a challenging yet worthwhile endeavor.
Asian Pop. Ratio European Pop. Ratio
Cantonese 87 0.78 English 1456 6.57
Mandarin 1138 3.80 German 133 18.41
Hindi 610 1.24 Russian 255 2.68
Arabic 376 4.82 French 310 7.34
Bengali 273 0.70 Italian 68 9.57
Persian 79 1.84 Spanish 559 2.54
Japanese 123 13.24 Polish 45 7.76
Marathi 99 1.14 Ukrainian 39 2.46
Turkish 90 4.16 Romanian 26 9.19
Korean 82 6.59 Dutch 55 8.56
Vietnamese 86 2.40 Swedish 11 35.55
Table 1: Language, population (Pop.), and publication to population ratio indirectly show the proportion of NLP resources to different languages (Appendix 7).
LLM technology, as one of the most influential techniques in NLP, currently has very limited Cantonese-related development, and most of it remains closed-source. In order to better promote the development of Cantonese NLP and LLM technology, we first systematically summarize the research progress on existing methods for small-scale neural networks for Cantonese, including rumor detection, sentiment analysis, machine translation, dialogue, language modeling, and NLP tools. Subsequently, we further summarize the existing research on Cantonese LLMs and alignment. Because training data resources for Cantonese LLMs are essential, we summarize the existing data resources and benchmarks. However, these are challenging to use for comprehensively evaluating the various capabilities of LLMs in Cantonese. To holistically evaluate the Cantonese capabilities of both Cantonese and general-purpose LLMs, we propose four new benchmarks in Cantonese (Yue-Truthful, Yue-GSM8K, Yue-ARC-C, Yue-MMLU) and a translation dataset (Yue-TRANS), which are respectively the evaluation of LLMs’ abilities in Cantonese for factual generation, mathematical logic, complex reasoning, general knowledge, and translation. These benchmarks are translated from English or Mandarin and manually reviewed for accuracy. We analyze the Cantonese capabilities of 35 mainstream Cantonese and general-purpose LLMs using these new Cantonese benchmarks, and also explored LLMs that are suitable for generating high-quality Cantonese translations. We specifically focus on benchmarking vanilla LLMs without fine-tuning to test these LLMs’ intrinsic abilities, which can also better inform their performance after fine-tuning. Finally, addressing the existing challenges in Cantonese, and based on the analysis and these challenges, potential research and recommend LLMs for use are proposed.
2 Cantonese existing NLP method
2.1 Cantonese small-scale neural network
Cantonese NLP based on small-scale neural network research encompasses a variety of domains such as rumor detection, sentiment analysis, machine translation, and dialogue, leveraging small neural network methods, models, and tools.
Rumor Detection. Chen et al. (2020) developed a dataset of 27,328 Cantonese tweets, divided into rumors and non-rumors, and introduced an attention-based model, XGA, which integrates XLNet and BiGRU to analyze semantic and sentiment aspects Chen et al. (2020); Yang et al. (2019). Chen et al. (2024) further developed CantoneseBERT to capture glyph and pronunciation clues of Cantonese characters, along with a Cantonese rumor detection model, SA-GCN, that uses the BiGCN model to encode global structural information of tweet hierarchies Chen et al. (2024).
Sentiment Analysis. Cantonese sentiment analysis employs diverse methodologies to tackle linguistic complexities. Early approaches used Naive Bayes and SVMs with character-based bi-grams, while later studies utilized Hidden Markov Models for text segmentation and part-of-speech tagging, developing emotion-specific dictionaries via rule-based systems Zhang et al. (2011); Chen et al. (2013, 2015). More recent studies have enhanced classification accuracy using both supervised and unsupervised methods across various domains, with Lee et al. (2019) exploring fine-grained emotion analysis across languages Ngai et al. (2018); Xiang et al. (2019); Lee (2019).
Machine Translation. Initial Cantonese machine translation research used heuristic rules and bilingual knowledge bases Zhang (1998); Wu et al. (2006), transitioning to statistical methods to address resource limitations Huang et al. (2016). Recent advancements include large-scale datasets and unsupervised models that utilize cross-lingual embeddings and Transformer architecture Liu (2022); Dare et al. (2023).
Dialogue Summarization and Generation. Lee et al. (2021) focused on generating questions and restating information in Cantonese dialogue systems, particularly enhancing performance in counseling chatbots by fine-tuning the BertSum model Lee et al. (2021); Liu and Lapata (2019). Lee also developed a dataset for virtual counselors to guide response selection through a regression model Lee and Liang (2021).
Cantonese Language Model. Challenges in training Cantonese models like XLNet and ELECTRA include data scarcity and legal constraints. Chen et al. (2024) introduced CantoneseBERT and the SA-GCN model for detailed analysis and rumor detection, utilizing permutation learning and adversarial training, though the training corpus included significant Standard Chinese content Chen et al. (2024); Yang et al. (2019); Clark et al. (2020).
Cantonese NLP Tools. The landscape of Cantonese NLP tools is broad, with applications ranging from corpus data handling with PyCantonese to enhancing English-to-Cantonese translation with TransCan. Tools like Cantonese Word Segmentation and cantoseg improve text accuracy, while canto-filter and songotsti support language identification Lee et al. (2022).
2.2 Cantonese large language model
Developing Cantonese LLMs faces challenges due to the unique linguistic features of Cantonese and limited data availability, necessitating comprehensive, high-quality datasets for effective pre-training. Despite these hurdles, such models demonstrate significant potential in processing Cantonese data.
there are very few large Cantonese models available, with Sensechat-5333https://www.sensetime.com/en/news-detail/51168164?categoryId=1072 being the only reliable non-commercial Cantonese LLM at present. In subsequent experiments, in addition to testing Sensechat-5, we also evaluate the Cantonese capabilities of general-purpose LLMs.
Recent research validates the effectiveness of ChatGPT in Cantonese dialogue and sentiment analysis, particularly in analyzing interactions from a Hong Kong web counseling service Fu et al. (2024). The introduction of the CanChat bot has improved emotional support for students in Hong Kong, particularly during and post the COVID-19 pandemic Fung et al. (2023).
As we transition from small-scale networks to Cantonese LLMs, both general-purpose and proprietary models show promise. However, quantifying their performance remains a challenge. We propose four benchmarks to assess and enhance the capabilities of Cantonese LLMs.
3 Cantonese data summary and new benchmarks construction
3.1 Existing Cantonese data
interactions, with Cantonese becoming the main focus of most bilingual dictionaries by the 19th century Xiang et al. (2024). Hong Kong led the development of Cantonese linguistic resources, including bilingual corpora from the Legislative Council Wu (1994), a one-million-character Cantonese corpus from children’s dialogues Hun-tak Lee (1999), and specialized corpora for Cantonese-speaking children Yip and Matthews (2007). Significant contributions also came from television and theater productions Leung and Law (2001), and the University of Hong Kong’s work on spontaneous speech, focusing on transcription and tagging Ping-Wai (2006). A parallel Cantonese-Standard Chinese corpus was developed for machine translation, sourced from television broadcasts Lee (2011). Recent efforts have focused on closing the data gap between Cantonese and other major languages through a small dependency treebank and a comprehensive bilingual dictionary, enhancing tools for translation Xiang et al. (2024).
3.2 New benchmarks construction
There are various benchmarks for testing the capabilities of LLMs, yet there are no publicly available benchmarks specifically designed to evaluate the proficiency of Cantonese LLMs. Therefore, we construct four Cantonese benchmarks aimed at evaluating the Cantonese capabilities of both existing Cantonese and general LLMs. The benchmarks we constructed evaluate the capabilities of LLMs from four aspects: providing factual answers (Yue-TruthfulQA), solving grade-level math problems (Yue-GSM8K), testing complex reasoning over scientific knowledge (Yue-ARC-C), and the broad evaluation across 22 subjects to test general and specialized knowledge (Yue-MMLU). The statistics of the datasets are as follows:
Datasets Number Types
Yue-TruthfulQA 817 Factual generation
Yue-GSM8K 1319 Mathematical logic
Yue-ARC-C 1171 Complex reasoning
Yue-MMLU 3721 General knowledge
Yue-TRANS 400 Translation
Table 2: Question number and type of the datasets.
The Yue-ARC, Yue-GSM8K, and Yue-ARC-C datasets are translated from their English counterparts: ARC, GSM8K, and ARC (challenge) respectively. The Yue-MMLU dataset is derived from CMMLU, featuring translations across an extensive range of twenty-two topics (Appendix A.6). Yue-TRANS consists of a randomly selected set of four hundred translation pairs [^4] https://huggingface.co/hon9kon9ize (two hundred pairs each from Mandarin to Cantonese and English to Cantonese).
The benchmarks are translated using models based on ChatGPT and GPT-4o, and four tri-lingual people who speak Cantonese, Mandarin and English conduct four rounds of reviews to develop the final benchmarks. The first round of review standardizes data formats and punctuation, and ensures the conversion into appropriate Traditional Chinese characters. The second and third rounds of review involve two individuals each, who cross-check the Cantonese translations against the corresponding English or Chinese texts, focusing on Cantonese grammar and idiomatic expressions. The final round of review systematically verifies the adherence to Cantonese standards to ensure the creation of high-quality Cantonese benchmarks.
Figure 3 shows that the first term "watermelon seeds" and the fourth term "color change" are colloquial expressions used in both everyday life and science in Cantonese. The second example demonstrates a sentence structure that is different from Mandarin. The third is a place name in Cantonese.
Figure 3: Translation example.
Models (Yue-TruthfulQA) 0-shot (correct) 5-shot (correct)
Rouge-l Bleu-4 BERTScore Rouge-l Bleu-4 BERTScore
Qwen-1.5-110b 26.04 15.95 69.29 31.73 19.53 70.87
Qwen-2-72b 10.86 9.68 65.62 17.52 12.38 67.72
Qwen-2.5-72b 13.03 9.64 66.94 20.23 12.87 69.53
Mixtral-8x22b 14.74 10.83 66.72 20.40 14.09 68.05
Mixtral-large-2 19.72 13.01 69.06 31.38 18.61 72.07
Llama-3-70b 10.98 9.51 66.10 33.06 19.31 71.95
Llama-3.1-70b 21.03 14.30 68.31 34.72 20.54 70.80
Phi-3-medium 18.70 12.00 67.36 22.00 13.72 67.57
Gemma-2-27b 8.09 8.44 64.41 11.33 9.98 63.66
Yi-1.5-34b 15.41 11.11 67.57 20.30 13.20 69.50
Internlm-2.5-20b-chat 6.96 7.73 62.99 3.28 6.06 66.99
Internlm-2.5-20b-turbomind 9.49 11.55 66.70 11.98 16.56 68.86
ERNIE-Turbo 17.91 11.30 66.71 21.19 12.19 68.29
Sensechat-5 24.75 15.11 68.43 32.45 19.70 70.02
Claude-3.5 14.23 9.95 67.56 12.66 10.06 68.12
GLM-4 13.44 10.07 67.26 23.57 14.28 70.30
ChatGPT 25.07 14.81 67.78 31.84 18.42 70.41
GPT-4o 17.58 12.17 68.68 27.64 16.52 71.59
GPT-4 19.47 13.45 68.99 28.43 16.74 71.26
Table 3: Results of the comparison between texts generated by various LLMs in Yue-TruthfulQA based on 0-shot and 5-shot settings and the correct texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text and semantics similarity (Table BBBBBB for more results).
Models Acc. (0-shot) Acc. (5-shot)
Qwen-1.5-110b 54.89 58.30
Qwen-2-72b 77.86 77.71
Qwen-2.5-72b 83.62 83.55
Mixtral-8x22b 65.20 66.19
Mixtral-large-2 80.14 81.27
Llama-3-70b 73.62 75.66
Llama-3.1-70b 53.60 79.00
Phi-3-medium 59.29 63.15
Gemma-2-27b 9.70 2.65
Yi-1.5-34b 69.45 69.45
Internlm-2.5-20b-chat 71.87 72.33
ERNIE-turbo 14.03 10.92
SenseChat-5 77.48 73.16
Claude-3.5 77.79 81.27
GLM-4 78.17 77.10
ChatGPT 23.35 41.09
GPT-4o 83.24 83.40
GPT-4 81.12 83.02
Table 4: Results of the comparison between various LLMs answer in Yue-GSM8K based on 0-shot and 5-shot and groundtruth (Table 918 for more results).
Models Acc. (0-shot) Acc. (5-shot)
Qwen-1.5-110b 88.64 90.09
Qwen-2-72b 88.64 88.56
Qwen-2.5-72b 92.74 92.91
Mixtral-8x22b 76.09 76.09
Mixtral-large-2 89.5 90.61
Llama-3-70b 85.06 84.97
Llama-3.1-70b 88.98 88.39
Phi-3-medium 77.63 78.31
Gemma-2-27b 67.98 55.59
Yi-1.5-34b 84.88 86.42
Internlm-2.5-20b-chat 82.15 82.58
ERNIE-turbo 44.41 46.46
SenseChat-5 88.47 87.28
Claude-3.5 91.55 92.23
GLM-4 88.9 88.73
ChatGPT 69.68 70.71
GPT-4o 91.97 94.45
GPT-4 92.66 92.06
Table 5: Results of the comparison between various LLMs answer in Yue-ARC-C based on 0-shot and 5-shot and groundtruth (Table 1019 for more results).
Models (Yue-MMLU) 0-shot (correct) 5-shot (correct)
STEM Hum. $S.$ C.S. Oth. STEM Hum. $S.$ C.S. Oth.
Qwen-1.5-110b 75.07 88.48 83.89 80.57 82.14 79.96 88.12 88.75 84.8 89.31
Qwen-2-72b 81.68 89.93 88.47 81.9 87.48 85.7 89.54 88.12 83.72 87.73
Qwen-2.5-72b 83.72 87.88 87.2 80.68 85.36 83.89 89.7 88.75 82.34 87.42
Mixtral-8x22b 50.4 57.08 59.28 44.02 48.76 58.94 59.72 62.44 49.78 57.83
Mixtral-large-2 60.38 76.08 74.92 60.19 70.74 68.5 79.65 78.84 63.85 71.66
Llama-3-70b 65.17 73.58 75.22 57.87 72.84 64.06 72.82 73.16 57.34 72.95
Llama-3.1-70b 67.32 76.57 76.93 60.96 73.56 72.23 78.13 78.23 64.16 74.9
Phi-3-medium 45.26 61.42 58.4 45.65 51.33 49.88 59.33 59.35 45.49 53.02
Gemma-2-27b 48.5 54.05 53.32 36.92 48.22 40.62 41.72 43.81 32.99 46.03
Yi-1.5-34b 68.48 81.92 81.74 70.89 79.76 74.13 85.12 83.38 78.2 80.3
Internlm-2.5-20b-chat 67.16 81.56 77.72 73.05 72.64 66.22 82.65 78.42 72.94 74.03
ERNIE-turbo 43.34 56.05 53.97 52.02 44.82 41.01 57.66 54.28 49.49 46.95
Sensechat-5 69.97 83.21 80.73 73.86 76.95 68.98 82 79.88 73.52 74.77
Claude-3.5 66.47 76.84 78.04 60.6 75.98 75.92 81.65 84.24 62.83 82.54
GLM-4 64.23 84.39 80.06 75.66 75.75 72.18 84.2 80.07 76 78.06
ChatGPT 49.78 58.13 58.74 45.46 52.42 60.28 59.81 60.61 47.5 54.54
GPT-4o 74.16 83.28 84.12 71.6 84.32 72.35 85.03 84.32 72.74 81.58
GPT-4 67.68 75.29 77.26 60.12 74.46 71.19 76.75 77.56 63.5 74.57
Table 6: Results of the comparison between texts generated by various LLMs in Yue-MMLU based on 0-shot and 5-shot settings and the correct texts (Table $B$ for more results).
4 Experiment and analysis
4.1 Implementation details
We conduct experiments on the Yue-ARC, Yue-MMLU, Yue-GSM8K, Yue-TruthfulQA, and Yue-TRNAS datasets. We use APIs and six A100-80G GPUs to perform inference with LLMs. We employ sampling hyperparameters with top-p set to 1.0 and a temperature of 0.2 for generation (Specific prompts in the Appendix A.9). We use xFinder Yu et al. (2024) to extract the answers of Yue-ARC-C, Yue-MMLU, Yue-GSM8K for later evaluation.
4.2 Evaluation
For Yue-TruthfulQA and Yue-TRANS (0-shot and 5-shot), we utilize Rouge-l, Bleu-4, and BERTScore as automatic evaluation metrics. Rouge-l Lin (2004) measures the longest common subsequence between generated and reference texts. Bleu-4 Papineni et al. (2002) evaluates n-gram overlap up to four words between generated and reference texts. BERTScore Zhang* et al. (2020) evaluates semantic similarity using BERT embeddings (we use bert-base-multilingual-cased [^5] https://huggingface.co/google-bert/bert-base-multilingual-cased for Cantonese evaluation and roberta-large [^6] https://huggingface.co/FacebookAI/roberta-large for English evaluation). For Yue-GSM8K, Yue-ARC-C, and Yue-MMLU (0-shot and 5-shot), we employ Accuracy (Acc.) as the evaluation metric.
4.3 Large language models for comparison
We evaluate the Cantonese abilities of 35 models, encompassing twelve series of open-source and closed-source general and Cantonese LLMs, across four benchmarks. The LLMs evaluated are as follows (Appendix A.7 for details): (1) Qwen series: Qwen-7b, Qwen-1.5-7b, Qwen-1.5-110b, Qwen-2-7b, Qwen-2-72b, Qwen-2.5-7b, Qwen-2.5-72b; (2) Mixtral series: Mixtral-8x22b, Mixtral-large-2; (3) Llama series: Llama-2-7b, Llama-3-8b, Llama-3-70b, Llama-3.1-8b, Llama-3.1-70b; (4) Phi series: Phi-3-medium; (5) Gemma series: Gemma-2-27b; (6) Yi series: Yi-6b, Yi-1.5-6b, Yi-1.5-34b; (7) Internlm series: Internlm-2-7b, Internlm-2-20b, Internlm-2.5-7b, Internlm-2.5-20b; (8) ERNIE series: ERNIE-Lite, ERNIE-Tiny, ERNIE-Speed, ERNIE-Turbo ; (9) Sensechat series: Sensechat-5 (Cantonese); (10) Claude series: Claude-3.5-sonnet; (11) GLM series: GLM-4; (12) GPT series: ChatGPT, GPT-4o, GPT-4.
4.4 Results and analysis
Figure 4: a, b, c, d represent the performance of various LLMs on Yue-TruthfulQA, Yue-GSM8K, Yue-ARC-C, and Yue-MMLU, in both 0-shot and 5-shot. e, f, g, h correspond to comparisons of performance between four benchmarks and their English or Mandarin version. i indicates the effectiveness of translating from Mandarin and English into Cantonese (Table 3BBBBBB4 ,918510193.2BB2122 for more results).
The performance of Cantonese LLMs still lags behind that in Mandarin and English, and 5-shot is better than 0-shot.
Rouge-l and Bleu-4 excel in evaluating the overlap between candidate and reference, making them suitable for key information extraction, outperforming metrics used in 0-shot and 5-shot (Figure 4a, b, c, d). The latter setting generally surpasses the former, illustrating the advantage of additional references in improving generation. Unlike these metrics, BERTScore excels in deep semantic evaluation, important for evaluating disparities in benchmarks between Cantonese and English. Mainstream LLMs perform better in English than in Cantonese (Figure 4e, f, g, h), highlighting their proficiency in widely used languages and relative under-performance in Cantonese (Table 3, Table B). Accuracy metrics in benchmarks with unique answers corroborate these findings (Table 4, Table 5, Table 3.2, Table 18, Table 19, Table B). 5-shot typically show higher accuracy than 0-shot (Figure 4a, b, c, d), and performance in mainstream languages like English and Mandarin surpasses that in Cantonese, emphasizing the need for more Cantonese-focused research and LLM development (Figure 4e, f, g, h).
Different series of models are suitable for various Cantonese tasks.
Qwen-1.5-110b and Mixtral-large-2 lead in Cantonese factual generation in 0-shot, and Llama-3/3.1-70b, GPT-series in 5-shot, surpassing Sensechat-5, Gemma-2-27b and Phi-3-medium, excluding smaller models, is prone to hallucinations, affecting its scores (Figure 4).
GPT-4, GPT-4o and Claude-3.5 excel in mathematical logic, followed by Mixtral-large-2, Llama-3.1-70b, and GLM-4. Models like ChatGPT perform better in English, indicating challenges in Cantonese mathematical reasoning due to language nuances (Table 4, Figure 4b, g).
For complex reasoning, GPT-4 and GPT-4o consistently demonstrates optimal performance, closely followed by Qwen-2.5-72b, Claude-3.5, and Mixtral-large-2, each of which also exhibits excellent performance (Table 5).
For tasks across various topics of the MMLU, Qwen-2.5-72b consistently exhibits the best performance (Table 3.2). We compile a table detailing the best models for various personas along with recommended open-source models (Figure 5).
Enhance training data volume by utilizing models that excel in translating Cantonese data.
High-quality Cantonese data is crucial for the pre-training or alignment of Cantonese LLMs, with translations from Standard Chinese proving more effective due to linguistic similarities (Figure 4i), as opposed to English (Table 2122). While models like Gemma-2-27b perform less effectively in English-to-Cantonese translation, closed-source models such as Sensechat-5 and GPT series show minimal quality difference between 0-shot and 5-shot settings. Prioritizing translations from Standard Chinese, then English, optimizes data quality. Regarding cost-effectiveness, using closed-source models like Sensechat-5-Cantonese, ChatGPT, and GPT-4o is advisable if API costs are negligible (Table 2122). Models like Mixtral-large-2, Llama-3.1-70b and Qwen-1.5-110b offer cost savings and high-quality translations in both settings (Table 24, Figure 4i). The Llama and Qwen series, while not the highest in output quality, provides the best speed and cost-effectiveness for translating datasets to Cantonese.
4.5 Case study
In addition to the results analyzed above, we find that Gemma-2-27b frequently encounters hallucination issues, which impair its ability to handle Cantonese tasks (Appendix C). Although Qwen-2-72b exhibits good performance, it sometimes outputs training data. Nonetheless, the Qwen series of models remains proficient in handling Cantonese tasks (Appendix C). Appendix C for more cases.
5 Challenges and opportunities
5.1 Existing Cantonese challenges
Colloquialism.
Cantonese differs significantly from Standard Chinese in its spoken vocabulary, posing unique challenges for NLP models initially trained on Mandarin Snow (2004); Xiang et al. (2024). These differences are particularly evident in informal settings such as speech transcription and online forums like Linkg, and Openrice. Although smaller compared to datasets for English and Standard Chinese models like BERTweet Nguyen et al. (2020) and MacBERT Cui et al. (2021), these platforms still provide a substantial text corpus for training Cantonese-specific models Hale (2001, 2016). The abundant unique expressions and slang in Cantonese, often embedded with complex cultural nuances, hinder adaptation of Standard Chinese-based models to Cantonese. For example, “Wan2 Sik6” literally means “looking for food”, but it is commonly used to describe seeking employment or earning money, carrying connotations of survival and making a living in Cantonese. In addition, common spelling mistakes and novel meanings in Cantonese further complicate model training, emphasizing the need for robust, Cantonese-specific vocabularies and corpora to capture the full breadth of colloquialisms and idioms of the language Li and Costa (2009).
Multilingualism.
To elucidate the multilingual dynamics in social media of Hong Kong, Xiang et al. (2024) identify frequent code-switching between Cantonese and Standard Chinese, and a significant presence of English Yue-Hashimoto (1991); Li (2006). Highlighting the multilingualism, examples include Cantonese sentences incorporating English terms, such as “deadline” seamlessly integrated as in “Gan2 M4 Cit3 deadline” (struggling to meet the deadline), and the use of the Japanese loanword “Kawaii” (cute), pronounced and adapted locally in phrases like “Ni1 Gin6 Saam1 Hou2 kawaii” (This shirt is very cute). These findings underscore the need for Cantonese NLP systems to handle multilingual code-switching and suggest adding spelling correction and dialect identification to improve data processing.
Figure 5: LLMs proficient in handling various tasks.
5.2 Opportunities
Given the existing challenges in Cantonese language and the evaluation results on benchmarks, we propose the following potential research directions and recommended models.
Data augmentation. Data augmentation methods for Cantonese are similar to those used broadly, including label-invariant methods that modify text while preserving labels Wei and Zou (2019); Min et al. (2020); Shi et al. (2021), and label-variant techniques that alter semantics for new instances Jin et al. (2019); Dai et al. (2019). Supervised contrastive learning enhances task-specific neural representations Sedghamiz et al. (2021), and LLM-based strategies are reviewed in Ding et al. (2024). For dataset conversion to Cantonese, high-capability models like Sensechat-5 and GPT-4 are recommended if costs allow (Table 2122, Table 24). Budget-friendly alternatives include Mixtral-large-2 and Llama-3.1-70b, with Llama models providing cost-effective speeds despite lower quality (Table 24).
Code-switch. Developments in LLMs suggest emergent abilities for untrained tasks, although effectiveness varies across scripts and languages Mann et al. (2020); Bang et al. (2023). Research in SCN-adapted LLMs is progressing, benefiting Cantonese NLP in the future Cui et al. (2023); Bai et al. (2023). We propose four benchmarks and compile a Yue-TRANS dataset, each involving two or more languages. Therefore, based on the performance observe on benchmarks, we recommend using newer versions of the Qwen, Llama, Mixtral, and Yi series (Figure 5).
Large language model. Based on the analysis above, we compile Figure 5, which presents the best LLMs in 0-shot and 5-shot, and suggests LLM series for various tasks. For work related to LLMs, we recommend using newer versions of the Qwen, Mixtral, Llama, and Yi series (Table B,910B). For tasks that involve only prompting, without the need for LLM training, we also recommend using closed-source models such as GPT, GLM, and Sensechat series models.
Limitations
The first limitation is the scarcity of work related to Cantonese LLMs, which restricts the extent of summarizing relevant studies. However, it is believed that with the publication of this paper, an increasing number of projects involving large-scale Cantonese models will be proposed. The second limitation is that the recommended LLMs presented in the article are for reference only; LLMs not recommended are not necessarily of inferior quality, nor does it imply they are unsuitable for Cantonese-related tasks. The selection of specific models for Cantonese-related tasks should be based on a detailed analysis of the specific issues at hand.
In addition, we specifically focus on benchmarking vanilla LLMs without fine-tuning to test these LLMs’ intrinsic abilities, which can also better inform their performance after fine-tuning.
Ethics Statement
Concerning the data annotators and the evaluation of data review, we ensure the selection of qualified tri-lingual individuals from Hong Kong and Guangdong who are compensated with reasonable hourly wages or other forms of subsidies as rewards. We have already obtained approval for this research from the Ethics Review Committee.
Appendix A Appendix
A.1 Cantonese speaking population statistics
Country/region Pop. Stat. Time
Hong Kong 6,529,000 2021
Macau 506,000 2016
Guangdong 40,000,000 2021
Guangxi 12,000,000 2022
Brunei 6,350 2006
Indonesia 180,000 1982
Malaysia 1,070,000 2000
Philippines 9,780 2000
Singapore 338,000 1993
Thailand 29,400 1984
Vietnam 862,000 1999
Table 7: Cantonese speaking population statistics. Pop. is population. Stat. Time is statistical time
A.2 Existing Cantonese data
At the end of the 16th century, Matteo Ricci compiles the first “Modern Bilingual Chinese Dictionary”, significantly incorporating Cantonese terms, highlighting its role in Sino-Western interactions. By the 19th century, most bilingual dictionaries focus on Cantonese Xiang et al. (2024). Historically, Hong Kong and related institutions lead Cantonese data initiatives. Wu (1994) creates a bilingual parallel corpus from the Hong Kong Legislative Council records, in both Standard Chinese and English. This effort is complemented by Hun-tak Lee (1999), who pioneers a Cantonese-only corpus with one million characters from dialogues involving children in Hong Kong, and Yip and Matthews (2007), who develops a bilingual corpus for Cantonese-speaking children. Additionally, a notable Cantonese corpus comes from television and theatrical productions in Hong Kong Leung and Law (2001). The University of Hong Kong further contributes by collecting and annotating spontaneous speech from dialogues and broadcasts, focusing on segmentation, parts of speech tagging, and phonetic transcription Ping-Wai (2006). Lee (2011) introduces a parallel corpus for machine translation between Cantonese and Standard Chinese, aligned at the sentence level, using data from Cantonese speeches on Hong Kong television and their Standard Chinese subtitles.
Recent efforts aim to bridge the data gap between Cantonese and other major languages. These include a small parallel dependency treebank for Cantonese and Mandarin, with 569 aligned sentences annotated using the Universal Dependencies scheme, and excerpts from the “ABC Cantonese-English Comprehensive Dictionary” providing 14,474 high-quality Cantonese-English parallel sentences, crucial for translation system development.
A.3 Cantonese small-scale neural network
Cantonese NLP research spreads across various topics, including rumor detection, sentiment analysis, machine translation, dialogue. We collect existing small neural network methods, models, and tools.
Rumor detection.
Chen et al. (2020) develop a dataset of 27,328 Cantonese tweets for rumor detection, split into 13,883 rumors and 13,445 non-rumors. They introduce an attention-based model, XGA, which combines XLNet Yang et al. (2019) and BiGRU to analyze both semantic and sentiment aspects. Chen et al. (2024) develop CantoneseBERT to capture glyph and pronunciation clues of Cantonese characters, and introduces a Cantonese rumor detection model, SA-GCN, that encodes global structural information of tweet hierarchies using the BiGCN model and extracts local semantic features with the CantoneseBERT model.
Sentiment analysis.
Cantonese sentiment analysis utilizes diverse methodologies to address its linguistic complexities. Zhang et al. (2011) apply Naive Bayes and SVMs with character-based bi-grams in the Openrice app for effective emotion detection. Chen et al. (2013, 2015) deploy Hidden Markov Models for text segmentation and part-of-speech tagging, developing emotion-specific dictionaries via rule-based systems. These studies demonstrate the value of combining machine learning with lexical techniques Zhang et al. (2011); Chen et al. (2013, 2015). In addition, Ngai et al. (2018) and Xiang et al. (2019) enhance classification accuracy using supervised and unsupervised methods in various domains. Lee (2019) explores fine-grained emotion analysis across languages, achieving significant results. These efforts underscore the importance of multi-methodological approaches Ngai et al. (2018); Xiang et al. (2019); Lee (2019). Tan et al. (2021) successfully employ Transformers pre-trained on simplified Chinese Tan et al. (2021).
Machine translation.
Initial research in this area utilizes heuristic rules, with significant contributions from Zhang (1998) and a bilingual Cantonese-English knowledge base by Wu et al. (2006). The focus has since shifted to statistical machine translation, exemplified by Huang et al. (2016), who addresses the challenges of translating between Cantonese and Mandarin with limited resources. Wong et al. (2018) improves this approach by enhancing parallel data for more efficient model training. Recent developments include a large-scale evaluation dataset by Liu (2022), containing over 35,000 Mandarin-Cantonese sentence pairs, and unsupervised translation models by Dare et al. (2023), which use cross-lingual embeddings and combine Transformer architecture with character-based tokenization to create a new corpus of approximately 1 million Cantonese sentences.
Dialogue summarization and generation.
Lee et al. (2021) explores generating questions and restating information in Cantonese dialogue systems, particularly for counseling chatbots. They enhance performance by fine-tuning the pre-trained BertSum model Liu and Lapata (2019) on Cantonese data, effective in tasks involving text summarization and question generation. In dialogue generation, Lee and Liang (2021) develops a specialized dataset for virtual counselors containing 1,028 post-reply pairs addressing test anxiety and loneliness, using these categories to guide response selection through a regression model.
Cantonese language model.
Training Cantonese language models like XLNet Yang et al. (2019) and ELECTRA Clark et al. (2020) from ToastyNews [^7] https://huggingface.co/toastynews faces challenges due to data scarcity and legal constraints. Chen et al. (2024) introduce CantoneseBERT and the SA-GCN model for detailed analysis and rumor detection in tweets, utilizing innovative methods like permutation learning and adversarial training. However, the training corpus largely includes Standard Chinese, leading to potential language contamination, and the impact on model efficacy remains unexplored.
Cantonese NLP tools.
The landscape of Cantonese NLP tools is diverse, addressing various needs. PyCantonese Lee et al. (2022) facilitates corpus data handling and linguistic analysis. Hong Kong Cantonese Localization provides culturally contextual translations. TransCan [^8] https://github.com/ayaka14732/TransCan enhances English-to-Cantonese translation, surpassing commercial solutions like Baidu and Bing. Text segmentation tools like Cantonese Word Segmentation [^9] https://github.com/wchan757/Cantonese_Word_Segmentation and cantoseg [^10] https://github.com/ayaka14732/cantoseg improve accuracy through custom dictionaries. canto-filter [^11] https://github.com/CanCLID/canto-filter categorizes texts based on linguistic features, while songotsti [^12] https://github.com/justinchuntingho/songotsti and fast-langid [^13] https://github.com/ffreemt/fast-langid offer additional support for language identification.
A.4 Cantonese large language model
Developing Cantonese LLMs is challenging due to scarce resources and the distinct features of the Cantonese language, necessitating extensive high-quality datasets for pre-training [^14] https://www.sensetime.com/en/news-detail/51168164?categoryId=1072. Despite these obstacles, these models show promising capabilities in processing Cantonese.
Aligning Cantonese LLMs for downstream tasks, such as prompting, supervised fine-tuning, and reinforcement learning from human feedback, is cost-effective and helps eliminate biases and meet cultural expectations.
Recent studies Fu et al. (2024) validate ChatGPT’s effectiveness in Cantonese dialogue and sentiment analysis, analyzing messages from a Hong Kong web counseling service. The CanChat bot, introduced to enhance counseling services in Hong Kong, provides initial support to students, improving their emotional well-being during and beyond the COVID-19 pandemic Fung et al. (2023).
Transitioning from small-scale networks to exploring Cantonese LLMs, both general-purpose and closed-source models show promise, but quantifying performance is challenging. We propose four benchmarks to evaluate and advance Cantonese capabilities in LLMs.
A.5 Evaluation tools
  • Rouge-l: from rouge_metric import PyRouge
  • Bleu-4: from nltk.translate.bleu_score import sentence_bleu, SmoothingFunction
  • BERTScore: bert-base-multilingual-cased & roberta-large
A.6 Yue-MMLU
We select twenty-two topics from CMMLU that cover most of the themes in CMMLU to serve as the topics for Yue-MMLU, which are as follows:
  • chinese_civil_service_exam
  • arts
  • electrical_engineering
  • chinese_literature
  • education
  • economics
  • ethnology
  • college_medicine
  • journalism
  • management
  • marketing
  • philosophy
  • security_study
  • sociology
  • world_history
  • world_religions
  • high_school_geography
  • machine_learning
  • marxist_theory
  • professional_psychology
  • sports_science
  • logical
A.7 Source of evaluation LLMs
This section covers the evaluation of LLMs along with the corresponding Hugging Face links and the names of the APIs.
A.8 Experimental results
A.8.1 Cantonese and English TruthfulQA (best and incorrect)
Table B (comparison between best answer and groundtruth) and Table B (comparison between incorrect answer and groundtruth) are the experimental results based on the Cantonese and English version of TruthfulQA.
A.8.2 English TruthfulQA (correct)
Table B (comparision vetween correct answer and groundtruth) is the experimental result based on the English version of TruthfulQA, intended for comparison with the Cantonese version of TruthfulQA. For more results, please refer to the publicly available evaluation platform [^15] https://huggingface.co/open-llm-leaderboard.
A.8.3 English GSM8K
Table 18 is the experimental result based on the English version of GSM8K, intended for comparison with the Cantonese version of GSM8K. For more results, please refer to the publicly available evaluation platform [^16] https://huggingface.co/open-llm-leaderboard.
A.8.4 English ARC challenge
Table 19 is the experimental result based on the English version of ARC Challenge, intended for comparison with the Cantonese version of ARC Challenge. For more results, please refer to the publicly available evaluation platform [^17] https://huggingface.co/open-llm-leaderboard.
A.8.5 CMMLU
Table B is the experimental result based on the Standard Chinese version of MMLU, intended for comparison with the Cantonese version of MMLU. For more results, please refer to the publicly available evaluation platform [^18] https://huggingface.co/open-llm-leaderboard.
A.8.6 Translation
Table 2122 is the experimental result based on the Yue-Trans datasets. Table 23 and Table 24 reflect the running time of different LLMs on the translation dataset.
A.9 Prompt templates for multilingual evaluation
This section details the prompt templates used for the Cantonese, English, and Standard Chinese datasets tested in our experiments. Each dataset was evaluated under both 0-shot and 5-shot settings. For the 5-shot setting, except for the translation task (Yue-TRANS), the prompts were generated using a sliding window approach, where the preceding five examples from the dataset (Yue-TruthfulQA, Yue-GSM8K, Yue-ARC-Challenge, and Yue-MMLU) were utilized as context for each new example. For the Yue-TRANS translation task, the BM25 algorithm was employed to identify and select the five most similar examples to serve as few-shot examples. Below, we outline the prompt structures and methodologies used for generating the few-shot examples.
The inference and evaluation processes in this study were facilitated by the OpenCompass platform, which provided a robust and universal evaluation framework for foundation models Contributors (2023).
{CJK}
UTF8bsmi
A.9.1 Yue-TruthfulQA prompt
0-shot:
    用粵語答下面問題:    問題:[QUESTION]    回應:
5-shot:
    樣例1-5:        問題:[EXAMPLE_QUESTION]        回應:[EXAMPLE_ANSWER]    用粵語答下面問題:    問題:[TARGET_QUESTION]    回應:
A.9.2 En-TruthfulQA prompt
0-shot:
    Answer the following question in English:    Question: [QUESTION]    Answer:
5-shot:
    Example 1-5:        Question: [EXAMPLE_QUESTION]        Answer: [EXAMPLE_ANSWER]    Answer the following question in English:    Question: [TARGET_QUESTION]    Answer:
{CJK}
UTF8bsmi
A.9.3 Yue-GSM8K prompt
0-shot:
    請逐步思考,最終答案前用「####」標記。用粵語答下面問題:    問題:[QUESTION]    用粵語回答問題:
5-shot:
    樣例1-5:        問題:[EXAMPLE_QUESTION]        回應:[EXAMPLE_ANSWER]    請逐步思考,最終答案前用「####」標記。用粵語答下面問題:    問題:[TARGET_QUESTION]    用粵語回答問題:
A.9.4 En-GSM8K prompt
0-shot:
    Please think step by step, mark the final answer with ’####’.    Answer the following question in English:    Question: [QUESTION]    Answer the question in English:
5-shot:
    Example 1-5:        Question: [EXAMPLE_QUESTION]        Response: [EXAMPLE_ANSWER]    Please think step by step, mark the final answer with ’####’.    Answer the following question in English:    Question: [TARGET_QUESTION]    Answer the question in English:
{CJK}
UTF8bsmi
A.9.5 Yue-ARC-C prompt
0-shot:
    問題:[QUESTION]    由提供嘅選項中直接用選項嘅字母作答。    回應:
5-shot:
    樣例1-5:        問題:[EXAMPLE_QUESTION]        回應:[EXAMPLE_ANSWER]    問題:[TARGET_QUESTION]    由提供嘅選項中直接用選項嘅字母作答。    回應:
A.9.6 En-ARC-C prompt
0-shot:
    Question: [QUESTION]    Answer with the option’s letter from the given choices directly.    Answer:
5-shot:
    Example 1-5:        Question: [EXAMPLE_QUESTION]        Answer: [EXAMPLE_ANSWER]    Question: [TARGET_QUESTION]    Answer with the option’s letter from the given choices directly.    Answer:
{CJK}
UTF8bsmi
A.9.7 Yue-MMLU prompt
0-shot:
    以下係關於[SUBJECT]嘅單項選擇題,請直接畀出正確答案嘅選項。    問題:[QUESTION]    答案:
5-shot:
    樣例1-5:        問題:[EXAMPLE_QUESTION]        回應:[EXAMPLE_ANSWER]    以下係關於[SUBJECT]嘅單項選擇題,請直接畀出正確答案嘅選項。    問題:[TARGET_QUESTION]    答案:
{CJK}
UTF8gbsn
A.9.8 Zh-CMMLU prompt
0-shot:
    以下是关于[SUBJECT]的单项选择题,请直接给出正确答案的选项。    题目:[QUESTION]    答案:
5-shot:
    样例1-5:        题目:[EXAMPLE_QUESTION]        答案:[EXAMPLE_ANSWER]    以下是关于[SUBJECT]的单项选择题,请直接给出正确答案的选项。    题目:[TARGET_QUESTION]    答案:
{CJK}
UTF8bsmi
A.9.9 Yue-TRANS prompt
0-shot:
    請將下面呢句/段話直接翻譯成粵語:[SOURCE_TEXT]
5-shot:
    樣例1-5:        請將下面呢句/段話直接翻譯成粵語:[EXAMPLE_SOURCE_TEXT]        翻譯:[EXAMPLE_TRANSLATION_TEXT]    根據上面嘅例子,請將下面呢句/段話直接翻譯成粵語:    [TARGET_SOURCE_TEXT]
Appendix B Result
Models (Yue-TruthfulQA) 0-shot (correct) 5-shot (correct)
Rouge-l Bleu-4 BERTScore Rouge-l Bleu-4 BERTScore
Qwen-7b 6.42 3.99 51.57 4.04 2.98 49.7
Qwen-1.5-7b 20.54 13.41 66.45 12.45 10.41 61.59
Qwen-1.5-110b 26.04 15.95 69.29 31.73 19.53 70.87
Qwen-2-7b 13.27 10.00 66.14 16.91 11.48 67.71
Qwen-2-72b 10.86 9.68 65.62 17.52 12.38 67.72
Qwen-2.5-7b 18.51 12.28 66.07 6.83 8.07 58.97
Qwen-2.5-72b 13.03 9.64 66.94 20.23 12.87 69.53
Mixtral-8x22b 14.74 10.83 66.72 20.40 14.09 68.05
Mixtral-large-2 19.72 13.01 69.06 31.38 18.61 72.07
Llama-2-7b 3.48 6.42 57.16 3.57 6.52 56.36
Llama-3-8b 8.40 8.68 64.37 28.68 16.43 70.82
Llama-3-70b 10.98 9.51 66.10 33.06 19.31 71.95
Llama-3.1-8b 13.82 10.33 66.97 26.18 15.20 70.28
Llama-3.1-70b 21.03 14.30 68.31 34.72 20.54 70.80
Phi-3-medium 18.70 12.00 67.36 22.00 13.72 67.57
Gemma-2-27b 8.09 8.44 64.41 11.33 9.98 63.66
Yi-6b 1.37 5.05 53.16 1.07 5.99 54.21
Yi-1.5-6b 1.21 4.60 42.15 1.04 6.15 53.85
Yi-1.5-34b 15.41 11.11 67.57 20.30 13.20 69.50
Internlm-7b 5.89 6.65 56.33 2.59 3.68 55.73
Internlm-7b-turbomind 5.91 6.71 56.71 2.77 3.82 55.57
Internlm-2-7b 7.93 10.21 63.81 17.66 16.62 33.33
Internlm-2-7b-chat 6.7 7.68 61.83 3.3 5.49 65.47
Internlm-2-7b-turbomind 8.09 10.53 64.3 17.69 16.99 63.68
Internlm-2.5-7b 8.96 10.53 66.11 10.3 14.47 67.73
Internlm-2.5-7b-chat 7.13 8 63.48 4.05 7.19 67.61
Internlm-2.5-7b-turbomind 8.93 10.46 65.75 10.12 14.39 67.14
Internlm-2.5-20b-chat 6.96 7.73 62.99 3.28 6.06 66.99
Internlm-2.5-20b-turbomind 9.49 11.55 66.70 11.98 16.56 68.86
ERNIE-Lite 20.58 12.23 67.64 20.69 12.27 68.45
ERNIE-Tiny 27.16 14.49 68.45 27.91 15.28 68.84
ERNIE-Speed 22.58 13.15 67.84 23.61 13.82 68.27
ERNIE-Turbo 17.91 11.30 66.71 21.19 12.19 68.29
Sensechat-5 24.75 15.11 68.43 32.45 19.70 70.02
Claude-3.5 14.23 9.95 67.56 12.66 10.06 68.12
GLM-4 13.44 10.07 67.26 23.57 14.28 70.30
ChatGPT 25.07 14.81 67.78 31.84 18.42 70.41
GPT-4o 17.58 12.17 68.68 27.64 16.52 71.59
GPT-4 19.47 13.45 68.99 28.43 16.74 71.26
Table 8: Results of the comparison between texts generated by various LLMs in Yue-TruthfulQA based on 0-shot and 5-shot settings and the correct texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text similarity.
Models Acc. (0-shot) Acc. (5-shot)
Qwen-7b 0.68 6.75
Qwen-1.5-7b 36.62 26.31
Qwen-1.5-110b 54.89 58.30
Qwen-2-7b 50.49 61.11
Qwen-2-72b 77.86 77.71
Qwen-2.5-7b 63.84 44.20
Qwen-2.5-72b 83.62 83.55
Mixtral-8x22b 65.20 66.19
Mixtral-large-2 80.14 81.27
Llama-2-7b 0.83 1.82
Llama-3-8b 52.46 49.66
Llama-3-70b 73.62 75.66
Llama-3.1-8b 63.91 61.64
Llama-3.1-70b 53.60 79.00
Phi-3-medium 59.29 63.15
Gemma-2-27b 9.70 2.65
Yi-6b 2.12 10.16
Yi-1.5-6b 3.94 3.49
Yi-1.5-34b 69.45 69.45
Internlm-7b-turbomind 4.55 9.48
Internlm-2-7b 11.90 22.21
Internlm-2-7b-chat 56.41 48.67
Internlm-2-7b-turbomind 11.37 23.96
Internlm-2-20b 12.81 8.87
Internlm-2-20b-chat 60.42 59.21
Internlm-2.5-7b 57.70 44.05
Internlm-2.5-7b-chat 65.96 64.67
Internlm-2.5-7b-turbomind 56.79 42.99
Internlm-2.5-20b-chat 71.87 72.33
Internlm-2.5-20b-turbomind 45.03 61.41
ERNIE-turbo 14.03 10.92
ERNIE-Speed 28.81 28.28
ERNIE-Lite 54.81 32.15
ERNIE-Tiny 2.73 3.94
SenseChat-5 77.48 73.16
Claude-3.5 77.79 81.27
GLM-4 78.17 77.10
ChatGPT 23.35 41.09
GPT-4o 83.24 83.40
GPT-4 81.12 83.02
Table 9: Results of the comparison between answer generated by various LLMs in Yue-GSM8K based on 0-shot and 5-shot settings and groundtruth.
Models Acc. (0-shot) Acc. (5-shot)
Qwen-7b 11.02 14.6
Qwen-1.5-7b 65.24 67.55
Qwen-1.5-110b 88.64 90.09
Qwen-2-7b 79.08 78.39
Qwen-2-72b 88.64 88.56
Qwen-2.5-7b 81.64 83.35
Qwen-2.5-72b 92.74 92.91
Mixtral-8x22b 76.09 76.09
Mixtral-large-2 89.5 90.61
Llama-2-7b 23.57 34.24
Llama-3-8b 70.11 53.8
Llama-3-70b 85.06 84.97
Llama-3.1-8b 69 67.81
Llama-3.1-70b 88.98 88.39
Phi-3-medium 77.63 78.31
Gemma-2-27b 67.98 55.59
Yi-6b 31 66.01
Yi-1.5-6b 34.59 66.7
Yi-1.5-34b 84.88 86.42
Internlm-7b-turbomind 44.75 55.34
Internlm-2-7b-turbomind 44.75 55.34
Internlm-2.5-7b 78.14 77.46
Internlm-2.5-7b-chat 81.21 79.85
Internlm-2.5-7b-turbomind 77.37 77.37
Internlm-2.5-20b-chat 82.15 82.58
Internlm-2.5-20b-turbomind 84.29 76.94
ERNIE-turbo 44.41 46.46
ERNIE-Speed 74.47 74.04
ERNIE-Lite 72.25 77.28
ERNIE-Tiny 34.67 32.88
SenseChat-5 88.47 87.28
Claude-3.5 91.55 92.23
GLM-4 88.9 88.73
ChatGPT 69.68 70.71
GPT-4o 91.97 94.45
GPT-4 92.66 92.06
Table 10: Results of the comparison between answer generated by various LLMs in Yue-ARC-C based on 0-shot and 5-shot settings and groundtruth.
Models (Yue-MMLU) 0-shot (correct) 5-shot (correct)
STEM Hum. $S.$ C.S. Oth. STEM Hum. $S.$ C.S. Oth.
Qwen-7b 10.1 12.95 12.12 11.61 7.96 9.98 15.96 14.48 13.33 13.26
Qwen-1.5-7b 46.28 61.65 56.57 50.02 53 60.14 70.09 65.55 58.31 65.02
Qwen-1.5-110b 75.07 88.48 83.89 80.57 82.14 79.96 88.12 88.75 84.8 89.31
Qwen-2-7b 70.06 81.04 80.07 69.54 76.04 74.08 80.45 80.7 73.7 79.52
Qwen-2-72b 81.68 89.93 88.47 81.9 87.48 85.7 89.54 88.12 83.72 87.73
Qwen-2.5-7b 72.86 81.66 78.25 66.56 75.19 78.05 80.37 78.99 69.82 78.86
Qwen-2.5-72b 83.72 87.88 87.2 80.68 85.36 83.89 89.7 88.75 82.34 87.42
Mixtral-8x22b 50.4 57.08 59.28 44.02 48.76 58.94 59.72 62.44 49.78 57.83
Mixtral-large-2 60.38 76.08 74.92 60.19 70.74 68.5 79.65 78.84 63.85 71.66
Llama-2-7b 23.34 23.84 23.76 22.78 24.52 27.48 30.4 31.76 28.9 24.38
Llama-3-8b 49.13 59.3 56.51 47.53 53.72 44.04 58.47 53.94 46.24 52.55
Llama-3-70b 65.17 73.58 75.22 57.87 72.84 64.06 72.82 73.16 57.34 72.95
Llama-3.1-8b 45.96 58.27 56.08 44.86 53.7 53.45 58.06 58.31 45.86 53.65
Llama-3.1-70b 67.32 76.57 76.93 60.96 73.56 72.23 78.13 78.23 64.16 74.9
Phi-3-medium 45.26 61.42 58.4 45.65 51.33 49.88 59.33 59.35 45.49 53.02
Gemma-2-27b 48.5 54.05 53.32 36.92 48.22 40.62 41.72 43.81 32.99 46.03
Yi-6b 36.46 67.62 57.32 57.42 50.06 58.11 72.14 68.4 60.56 68.46
Yi-1.5-6b 17.34 35.98 38.77 32.9 25 58.53 67.89 66.56 60 62.05
Yi-1.5-34b 68.48 81.92 81.74 70.89 79.76 74.13 85.12 83.38 78.2 80.3
Internlm-7b-turbomind 31.9 48.79 44.03 41.14 39.82 39.84 51.74 50.06 43.6 42.32
Internlm-2-7b-turbomind 51.69 70.92 64.71 59.31 58.93 53.11 68.51 62.68 59.77 58.14
Internlm-2.5-7b 65.34 82.43 79.24 73.11 74.15 66.73 81.06 77.8 71.65 75.37
Internlm-2.5-7b-chat 64.4 80.92 76.8 70.24 75.02 65.04 80.84 76.79 70.47 75.19
Internlm-2.5-7b-turbomind 65.34 82.43 79.24 73.11 74.15 66.73 81.06 77.8 71.65 75.37
Internlm-2.5-20b-chat 67.16 81.56 77.72 73.05 72.64 66.22 82.65 78.42 72.94 74.03
Internlm-2.5-20b-turbomind 72.86 86.1 82.14 79.06 74.7 69.65 78.79 76.56 70.28 77.2
ERNIE-Lite 53.45 67.56 67.73 61.21 61.21 60.74 70.27 71.5 62.43 64.84
ERNIE-Tiny 34.78 37.86 37.88 33.08 32.29 32.52 38.63 37.58 32.52 34.6
ERNIE-turbo 43.34 56.05 53.97 52.02 44.82 41.01 57.66 54.28 49.49 46.95
Sensechat-5 69.97 83.21 80.73 73.86 76.95 68.98 82 79.88 73.52 74.77
Claude-3.5 66.47 76.84 78.04 60.6 75.98 75.92 81.65 84.24 62.83 82.54
GLM-4 64.23 84.39 80.06 75.66 75.75 72.18 84.2 80.07 76 78.06
ChatGPT 49.78 58.13 58.74 45.46 52.42 60.28 59.81 60.61 47.5 54.54
GPT-4o 74.16 83.28 84.12 71.6 84.32 72.35 85.03 84.32 72.74 81.58
GPT-4 67.68 75.29 77.26 60.12 74.46 71.19 76.75 77.56 63.5 74.57
Table 11: Results of the comparison between texts generated by various LLMs in Yue-MMLU based on 0-shot and 5-shot settings and the correct texts.
Models Mode Huggingface link & API name
Qwen-7b Huggingface https://huggingface.co/Qwen/Qwen-7B
Qwen-1.5-7b Huggingface https://huggingface.co/Qwen/Qwen1.5-7B
Qwen-1.5-110b Huggingface https://huggingface.co/Qwen/Qwen1.5-110B
Qwen-2-7b Huggingface https://huggingface.co/Qwen/Qwen2-7B-Instruct
Qwen-2-72b Huggingface https://huggingface.co/Qwen/Qwen2-72B-Instruct
Qwen-2.5-7b Huggingface https://huggingface.co/Qwen/Qwen2.5-7B
Qwen-2.5-72b Huggingface https://huggingface.co/Qwen/Qwen2.5-72B
Mixtral-8x22b Huggingface https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1
Mixtral-large-2 Huggingface https://huggingface.co/mistralai/Mistral-Large-Instruct-2407
Llama-2-7b Huggingface https://huggingface.co/meta-llama/Llama-2-7b
Llama-3-8b Huggingface https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
Llama-3-70b Huggingface https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct
Llama-3.1-8b Huggingface https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct
Llama-3.1-70b Huggingface https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct
Phi-3-medium Huggingface https://huggingface.co/microsoft/Phi-3-medium-128k-instruct
Gemma-2-27b Huggingface https://huggingface.co/google/gemma-2-27b-it
Yi-6b Huggingface https://huggingface.co/01-ai/Yi-6B
Yi-1.5-6b Huggingface https://huggingface.co/01-ai/Yi-1.5-6B-Chat
Yi-1.5-34b Huggingface https://huggingface.co/01-ai/Yi-1.5-34B-Chat
ERNIE-turbo API API: ERNIE-Bot-turbo
ERNIE-Speed API API: ERNIE-Speed-128K
ERNIE-Lite API API: ERNIE-Lite-8K
ERNIE-Tiny API API: ERNIE-Tiny-8K
Internlm-2-7b Huggingface https://huggingface.co/internlm/internlm2-7b
Internlm-2-7b-chat Huggingface https://huggingface.co/internlm/internlm2-20b
Internlm-2-20b-chat Huggingface https://huggingface.co/internlm/internlm2-chat-20b
Internlm-2.5-7b Huggingface https://huggingface.co/internlm/internlm2_5-7b
Internlm-2.5-7b-chat Huggingface https://huggingface.co/internlm/internlm2_5-7b-chat
Internlm-2.5-20b-chat Huggingface https://huggingface.co/internlm/internlm2_5-20b-chat
SenseChat-5 API API: SenseChat-5-Cantonese
Claude-3.5 API API: claude-3.5-sonnot-20240620
GLM-4 API API: GLM-4-0520
ChatGPT API API: gpt-3.5-turbo-instruct & gpt-3.5-turbo
GPT-4o API API: gpt-4o
GPT-4 API API: gpt-4-0125-preview
Table 12: The mode of the evaluation LLMs and their corresponding huggingface links & names of APIs.
Models (Yue-TruthfulQA) 0-shot (best) 5-shot (best)
Bleu-4 BERTScore Bleu-4 BERTScore
Qwen-7b 3.01 51.03 2.19 48.82
Qwen-1.5-7b 9.8 65.65 8.19 59.48
Qwen-1.5-110b 11.17 69.14 14.22 73.40
Qwen-2-7b 8.00 64.11 9.09 66.41
Qwen-2-72b 7.77 62.22 9.99 65.32
Qwen-2.5-7b 8.99 65.58 6.98 54.62
Qwen-2.5-72b 7.8 64.10 10.18 68.19
Mixtral-8x22b 8.54 64.63 11.31 67.43
Mixtral-large-2 10.01 67.37 14.14 73.41
Llama-2-7b 5.36 52.10 5.53 51.12
Llama-3-8b 7.26 60.79 12.94 71.77
Llama-3-70b 7.70 63.08 14.68 73.97
Llama-3.1-8b 8.19 63.97 11.93 70.64
Llama-3.1-70b 10.42 67.19 15.36 75.80
Phi-3-medium 9.34 65.84 10.98 66.81
Gemma-2-27b 7.15 60.94 8.14 61.54
Yi-6b 3.95 49.13 4.98 48.93
Yi-1.5-6b 3.82 38.22 5.15 48.43
Yi-1.5-34b 8.80 65.25 10.55 67.88
Internlm-7b 5.39 52.10 5.42 50.65
Internlm-7b-turbomind 5.26 52.45 5.4 50.56
Internlm-2-7b 6.51 64.20 5.58 33.94
Internlm-2-7b-chat 6.4 58.87 9.41 65.32
Internlm-2-7b-turbomind 6.85 64.77 9.71 64.19
Internlm-2-20b 8.65 68.08 2.61 20.27
Internlm-2-20b-chat 6.08 56.94 10.57 66.29
Internlm-2.5-7b 8.24 63.72 11.02 67.35
Internlm-2.5-7b-chat 6.79 60.35 8.41 65.13
Internlm-2.5-7b-turbomind 8.27 63.25 10.6 66.58
Internlm-2.5-20b-chat 6.55 59.61 8.06 64.32
Internlm-2.5-20b-turbomind 8.75 65.18 11.97 69.30
ERNIE-Lite 9.05 67.61 9.44 67.68
ERNIE-Tiny 14.49 70.05 10.82 70.39
ERNIE-Speed 9.54 68.33 10.49 68.49
ERNIE-Turbo 9.04 65.20 9.66 67.39
Sensechat-5 10.47 68.93 14.51 73.38
Claude-3.5 7.95 64.83 8.24 64.84
GLM-4 7.92 64.28 11.11 69.65
ChatGPT 10.42 67.84 13.82 71.87
GPT-4o 9.34 66.25 12.61 71.51
GPT-4 9.97 67.08 12.87 72.00
Table 13: Results of the comparison between texts generated by various LLMs in Cantonese version of TruthfulQA based on 0-shot and 5-shot settings and the best texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text similarity.
Models (TruthfulQA-English) 0-shot (best) 5-shot (best)
Bleu-4 BERTScore Bleu-4 BERTScore
Qwen-1.5-110b 12.78 85.83 20.10 87.19
Qwen-2-7b 8.76 83.80 16.37 87.10
Qwen-2-72b 6.99 81.36 8.58 82.97
Qwen-2.5-72b 9.22 84.30 11.33 85.72
Mixtral-8x22b 10.82 85.68 17.65 88.24
Mixtral-large-2 11.95 85.68 25.12 89.97
Llama-3-8b 10.04 83.86 32.17 90.98
Llama-3-70b 9.07 83.42 31.85 90.99
Llama-3.1-8b 9.81 83.19 31.18 90.56
Llama-3.1-70b 11.27 84.01 35.02 91.60
Phi-3-medium 12.33 86.70 24.27 89.57
Gemma-2-27b 8.46 83.20 10.52 84.24
Yi-1.5-34b 11.01 84.72 22.50 88.79
Internlm-2-7b 22.39 88.41 25.76 67.10
Internlm-2-7b-chat 8.41 83.21 16.14 86.96
Internlm-2-20b 21.77 88.38 26.70 86.60
Internlm-2-20b-chat 7.32 81.76 20.57 87.38
Internlm-2.5-7b 15.17 86.40 22.06 88.43
Internlm-2.5-7b-chat 7.77 82.73 9.95 84.40
ChatGPT 17.97 87.65 26.69 90.27
GPT-4o 10.93 85.28 32.38 90.94
GPT-4 11.51 85.16 34.34 91.36
Table 14: Results of the comparison between texts generated by various LLMs in English version of TruthfulQA based on 0-shot and 5-shot settings and the best texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text similarity.
Models (Yue-TruthfulQA) 0-shot (incorrect) 5-shot (incorrect)
Bleu-4 BERTScore Bleu-4 BERTScore
Qwen-7b 3.22 52.45 2.38 50.82
Qwen-1.5-7b 11.39 66.76 8.43 61.74
Qwen-1.5-110b 12.83 69.22 12.67 68.67
Qwen-2-7b 8.38 65.10 8.38 65.56
Qwen-2-72b 8.15 64.44 9.17 66.03
Qwen-2.5-7b 10.14 66.13 7.10 59.77
Qwen-2.5-72b 8.19 65.49 9.82 67.49
Mixtral-8x22b 9.24 66.27 10.14 66.11
Mixtral-large-2 10.60 68.40 12.62 69.74
Llama-2-7b 5.74 59.48 5.69 58.38
Llama-3-8b 7.69 64.07 11.03 68.54
Llama-3-70b 8.12 65.49 12.11 69.10
Llama-3.1-8b 8.72 66.38 10.73 68.22
Llama-3.1-70b 10.79 67.80 12.38 68.28
Phi-3-medium 10.23 67.07 10.40 66.07
Gemma-2-27b 7.40 63.04 8.05 62.28
Yi-6b 4.27 54.49 5.29 55.44
Yi-1.5-6b 4.15 43.31 5.35 55.07
Yi-1.5-34b 9.16 66.67 10.04 67.68
Internlm-7b 5.89 57.93 5.58 56.81
Internlm-7b-turbomind 5.91 58.23 5.54 56.7
Internlm-2-7b 7.93 64.39 4.73 32.66
Internlm-2-7b-chat 6.70 61.13 9.17 64.11
Internlm-2-7b-turbomind 8.09 64.76 8.50 62.9
Internlm-2-20b 10.24 66.74 2.30 21.15
Internlm-2-20b-chat 6.27 59.46 9.56 64.82
Internlm-2.5-7b 8.96 65.89 10.25 66.48
Internlm-2.5-7b-chat 7.13 62.94 8.84 66.68
Internlm-2.5-7b-turbomind 8.93 65.7 9.81 66.14
Internlm-2.5-20b-chat 6.96 62.15 8.23 65.67
Internlm-2.5-20b-turbomind 9.49 66.25 11.45 67.84
ERNIE-Lite 9.72 66.86 9.40 66.73
ERNIE-Tiny 11.50 67.96 11.63 67.90
ERNIE-Speed 10.18 66.93 10.52 66.93
ERNIE-Turbo 9.52 66.15 9.70 66.76
Sensechat-5 12.02 68.33 12.31 67.80
Claude-3.5 8.20 65.93 7.78 65.57
GLM-4 8.43 66.00 10.34 68.09
ChatGPT 11.29 67.46 13.07 68.69
GPT-4o 9.64 67.40 11.21 68.89
GPT-4 10.45 67.72 11.49 68.52
Table 15: Results of the comparison between texts generated by various LLMs in Cantonese version of TruthfulQA based on 0-shot and 5-shot settings and the incorrect texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text similarity.
Models (TruthfulQA-English) 0-shot (incorrect) 5-shot (incorrect)
Bleu-4 BERTScore Bleu-4 BERTScore
Qwen-1.5-110b 12.83 85.75 13.89 85.31
Qwen-2-7b 8.65 83.70 11.39 85.02
Qwen-2-72b 6.84 81.59 7.98 82.30
Qwen-2.5-72b 6.84 84.04 7.98 85.19
Mixtral-8x22b 9.94 85.19 12.63 86.15
Mixtral-large-2 11.18 85.21 16.21 86.50
Llama-3-8b 10.01 84.02 19.84 86.68
Llama-3-70b 8.68 83.55 18.89 86.80
Llama-3.1-8b 9.65 83.36 19.26 86.70
Llama-3.1-70b 10.86 83.95 19.27 86.64
Phi-3-medium 13.45 86.14 16.37 86.76
Gemma-2-27b 8.08 83.05 9.24 83.61
Yi-1.5-34b 10.63 84.48 15.49 86.31
Internlm-2-7b 23.38 87.47 17.54 64.53
Internlm-2-7b-chat 8.45 83.39 12.24 85.28
Internlm-2-20b 22.13 87.69 20.50 84.8
Internlm-2-20b-chat 7.20 81.94 14.08 84.78
Internlm-2.5-7b 15.76 86.17 16.10 86.39
Internlm-2.5-7b-chat 7.79 82.87 9.05 84.08
ChatGPT 17.78 87.22 20.45 87.50
GPT-4o 9.99 84.72 18.70 86.73
GPT-4 10.72 84.87 19.54 86.53
Table 16: Results of the comparison between texts generated by various LLMs in English version of TruthfulQA based on 0-shot and 5-shot settings and the incorrect texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text similarity.
Models (English-TruthfulQA) 0-shot (correct) 5-shot (correct)
Rouge-l Bleu-4 BERTScore Rouge-l Bleu-4 BERTScore
Qwen-1.5-110b 22.57 15.54 85.78 29.44 23.14 86.35
Qwen-2-7b 10.98 10.20 83.86 23.67 18.60 86.09
Qwen-2-72b 3.03 7.58 81.78 7.45 9.59 82.98
Qwen-2.5-72b 13.05 10.83 84.5 21.16 13.65 85.71
Mixtral-8x22b 18.59 12.91 85.78 31.05 20.61 87.58
Mixtral-large-2 20.57 14.63 85.69 41.46 28.92 88.30
Llama-3-8b 16.89 11.59 84.11 58.34 38.35 88.50
Llama-3-70b 12.09 10.46 83.84 53.00 36.77 88.94
Llama-3.1-8b 14.13 11.34 83.46 51.70 36.95 88.47
Llama-3.1-70b 18.12 13.24 84.18 55.22 40.54 88.88
Phi-3-medium 27.90 17.35 86.48 43.02 28.62 88.24
Gemma-2-27b 12.31 9.84 83.56 18.25 12.25 84.31
Yi-1.5-34b 17.22 13.22 84.79 35.33 25.82 87.56
Internlm-2-7b 47.58 28.78 87.13 41.57 30.32 65.51
Internlm-2-7b-chat 9.54 9.69 83.42 23.39 18.97 86.29
Internlm-2-20b 43.50 27.33 87.5 41.13 31.64 85.39
Internlm-2-20b-chat 4.81 8.14 82.11 31.44 24.45 85.8
Internlm-2.5-7b 34.44 18.62 86.06 39.19 25.39 87.31
Internlm-2.5-7b-chat 7.45 8.82 82.92 12.92 11.29 84.39
ChatGPT 37.81 21.95 87.20 50.43 31.44 88.55
GPT-4o 17.93 13.05 85.38 49.52 37.44 88.62
GPT-4 19.58 14.10 85.19 53.18 39.22 88.85
Table 17: Results of the comparison between texts generated by various LLMs in English-TruthfulQA based on 0-shot and 5-shot settings and the correct texts. Rouge-l, Bleu-4, and BERTScore are evaluation metrics for comparing text similarity.
Models Acc. (0-shot) Acc. (5-shot)
Qwen-1.5-110b 88.55 88.93
Qwen-2-7b 84.15 84.76
Qwen-2-72b 92.8 91.58
Qwen-2.5-72b 93.25 96.13
Mixtral-8x22b 91.51 91.58
Mixtral-large-2 95.38 95.15
Llama-3-8b 80.36 81.05
Llama-3-70b 93.4 93.33
Llama-3.1-8b 85.97 86.35
Llama-3.1-70b 95.3 95.3
Phi-3-medium 90.3 90.83
Gemma-2-27b 24.49 9.86
Yi-1.5-34b 87.95 88.4
Internlm-2-7b 46.63 61.56
Internlm-2-7b-chat 73.54 66.64
Internlm-2-20b 78.54 64.14
Internlm-2-20b-chat 78.54 75.28
Internlm-2.5-7b 77.48 65.88
Internlm-2.5-7b-chat 84.99 82.71
ChatGPT 65.28 67.25
GPT-4o 95.22 95.68
GPT-4 95 94.77
Table 18: Results of the comparison between answer generated by various LLMs in English-GSM8K based on 0-shot and 5-shot settings and groundtruth.
Models Acc. (0-shot) Acc. (5-shot)
Qwen-1.5-110b 82.66 77.6
Qwen-2-7b 65.41 69.7
Qwen-2-72b 69.79 79.83
Qwen-2.5-72b 95.19 94.76
Mixtral-8x22b 90.82 88.07
Mixtral-large-2 94.51 94.59
Llama-3-8b 81.63 78.88
Llama-3-70b 93.22 92.62
Llama-3.1-8b 80.52 84.21
Llama-3.1-70b 93.56 93.3
Phi-3-medium 93.13 92.1
Gemma-2-27b 82.92 72.79
Yi-1.5-34b 92.36 92.53
Internlm-2.5-7b 85.58 85.15
Internlm-2.5-7b-chat 87.04 86.78
Table 19: Results of the comparison between answer generated by various LLMs in English-ARC challenge based on 0-shot and 5-shot settings and groundtruth.
Models (Standard Chinese-MMLU) 0-shot (correct) 5-shot (correct)
STEM Hum. $S.$ C.S. Oth. STEM Hum. $S.$ C.S. Oth.
Qwen-1.5-110b 78.06 87.6 85.88 81.83 84.04 85.1 90.77 91.07 85.84 91.56
Qwen-2-7b 77.52 86.63 85.1 77.37 83.41 81.62 86.94 85.09 80.06 83.84
Qwen-2-72b 83.36 89.69 88.75 83.16 86.58 90.07 93.18 92.97 88.64 91.07
Qwen-2.5-72b 83.26 89.54 89.14 82.04 88.33 85.87 90.6 90.25 84.15 88.4
Mixtral-8x22b 57.88 63.27 64.51 49.18 57.28 62.38 62.97 63.7 51.52 58.26
Mixtral-large-2 68.49 79.48 77.03 64.36 70.8 71.65 81.95 78.76 66.87 74.52
Llama-3-8b 54.04 61.35 59.17 45.67 56.28 47.66 59.26 58 44.72 53.54
Llama-3-70b 72.64 77.23 77.44 60.22 76.3 72.04 75.31 74.99 58.74 74.72
Llama-3.1-8b 49.08 61.05 59.17 44.15 53.11 55.62 62.58 61.02 46.43 56.27
llama-3.1-70b 69.84 77.77 76.9 62.34 75.02 72.4 77.95 78.57 61.6 75.75
Phi-3-medium 58.54 63.46 65.61 48.45 61.5 57.18 62.84 66.32 49.76 59.06
Gemma2-27b 49.67 53.63 57.23 42.36 50.35 40.25 43.15 47.77 37.14 46.34
Yi-1.5-34b 73.02 83.78 82.99 74.6 83.72 78.87 86.24 84.47 77.68 85.06
Internlm-2.5-7b 75.62 88 83.95 79.14 80.86 70.52 87.27 83.38 79.6 80.19
Internlm-2.5-7b-chat 73.04 87.42 84.23 77.62 85.29 69.24 86.45 83.78 77.93 83.46
Table 20: Results of the comparison between texts generated by various LLMs in CMMLU based on 0-shot and 5-shot settings and the correct texts.
Models (mdn-yue) 0-shot 5-shot
Rouge-l Bleu-4 BERTScore Rouge-l Bleu-4 BERTScore
Qwen-7b 8.49 5.03 43.76 18.55 14.26 54.19
Qwen-1.5-7b 30.81 17.54 66.88 33.84 27.14 71.32
Qwen-1.5-110b 30.03 22.88 51.94 88.72 79.60 94.34
Qwen-2-7b 47.06 25.16 75.43 69.86 50.14 84.32
Qwen-2-72b 24.54 19.74 68.85 9.96 11.08 64.23
Qwen-2.5-7b 11.65 8.82 53.61 67.38 49.26 84.16
Qwen-2.5-72b 85.11 61.81 91.78 87.9 69.39 93.28
Mixtral-8x22b 46.7 32.04 74.81 65.75 51.59 84.47
Mixtral-large-2 85.71 64.83 91.99 88.55 72.7 93.42
Llama-2-7b 12.96 7.42 53.60 28.63 15.07 66.35
Llama-3-8b 26.69 33.14 74.81 56.04 43.53 84.47
Llama-3-70b 27.12 37.77 73.91 59.36 60.16 85.17
Llama-3.1-8b 69.88 44.3 84.67 82.33 61.34 90.39
Llama-3.1-70b 85.05 63.23 91.86 89.8 76.17 94.45
Phi-3-medium 66.73 36.79 83.65 76.53 48.58 88.49
Gemma-2-27b 9.16 11.3 62.11 7.39 8.56 59.14
Yi-6b 4.54 6.92 61.09 12.64 12.05 64.04
Yi-1.5-6b 7.3 8.29 63.56 23.01 19.81 68.54
Yi-1.5-34b 75.46 47.27 89.93 85.69 66.99 91.10
Internlm-7b-turbomind 4.26 6.42 58.07 13.33 12.25 64.46
Internlm-2-7b-turbomind 49.33 18.42 79.39 66.45 36.3 84.16
Internlm-2.5-7b 51 22.25 81.18 67.2 41.78 79.57
Internlm-2.5-7b-chat 47.97 16.95 81.13 66.38 34.87 86.43
Internlm-2.5-7b-turbomind 48.45 20.89 79.70 71.1 44.14 85.90
Internlm-2.5-20b-chat 36.62 23.41 74.42 65.29 43.79 82.97
Internlm-2.5-20b-turbomind 65.86 46.32 83.63 76.44 60.66 87.85
Sensechat-5 88.92 72.78 94.00 90.94 77.65 95.05
GLM-4 82.82 59.53 89.67 84.26 64.87 89.83
ChatGPT 86.47 68.02 92.09 87.46 73.62 91.49
GPT-4o 89.69 73.7 93.34 91.16 79.06 94.21
GPT-4 87.11 68.25 92.52 89.24 75.65 93.92
Table 21: Result based on the Yue-Trans datasets (translated from Mandarin to Cantonese).
Models (en-yue) 0-shot 5-shot
Rouge-l Bleu-4 BERTScore Rouge-l Bleu-4 BERTScore
Qwen-7b 1.72 0.61 43.76 15.11 7.06 54.19
Qwen-1.5-7b 27.55 9.08 66.88 58.75 27.57 71.32
Qwen-1.5-110b 2.75 1.09 51.94 74.39 40.05 94.34
Qwen-2-7b 50.85 21.26 75.43 68.58 31.62 84.32
Qwen-2-72b 34.17 19.05 68.85 14.4 14.58 64.23
Qwen-2.5-7b 7.91 5.07 53.61 39.88 21.58 84.16
Qwen-2.5-72b 70.95 33.72 91.78 73.36 37.8 93.28
Mixtral-8x22b 51.52 18.42 74.81 68.73 31.15 84.47
Mixtral-large-2 69.15 31.18 91.99 74.11 38.97 93.42
Llama-2-7b 0.91 2.21 53.60 21.47 11.09 66.35
Llama-3-8b 36.56 21.68 74.81 64.3 30.19 84.47
Llama-3-70b 58.58 28.11 73.91 61.72 34.58 85.17
Llama-3.1-8b 62.44 25.25 84.67 68.54 31.99 90.39
Llama-3.1-70b 66.05 29.71 91.86 73.38 37.78 94.45
Phi-3-medium 49.78 15.94 83.65 61.71 24.66 88.49
Gemma-2-27b 14.57 12.52 62.11 7.54 8.69 59.14
Yi-6b 4.91 4.27 61.09 15.98 14.51 64.04
Yi-1.5-6b 6.63 3.16 63.56 30.44 19.31 68.54
Yi-1.5-34b 65.15 27.91 89.93 71.36 35.06 91.10
Internlm-7b-turbomind 4.68 3.19 58.07 21.75 15.26 64.46
Internlm-2-7b-turbomind 34.2 10.72 79.39 57.74 24.81 84.16
Internlm-2.5-7b 29.76 12.15 81.18 14.38 6.04 79.57
Internlm-2.5-7b-chat 44.63 14.02 81.13 65.93 29.61 86.43
Internlm-2.5-7b-turbomind 29.84 12.29 79.70 38.72 16.61 85.90
Internlm-2.5-20b-chat 52.19 21.12 74.42 68.31 33.13 82.97
Internlm-2.5-20b-turbomind 48.89 17.96 83.63 68.51 35.97 87.85
Sensechat-5 66.95 33.91 94.00 74.02 39.04 95.05
GLM-4 70.73 34.26 89.67 72.93 38.07 89.83
ChatGPT 70.78 33.2 92.09 73.02 36.78 91.49
GPT-4o 72.84 36.34 93.34 74.4 39.85 94.21
GPT-4 72.01 34.42 92.52 73.89 37.38 93.92
Table 22: Result based on the Yue-Trans datasets (translated from English to Cantonese).
Models Total running time Number of GPU Batch size
Qwen-1.5-110b 11053.46 6 4
Qwen-2-7b 1463.17 1 8
Qwen-2-72b 21467.50 6 8
Mixtral-8x22b 19345.82 6 4
Mixtral-large-2 12302.97 6 4
Llama-3-8b 1449.98 1 8
Llama-3-70b 3741.66 6 16
Llama-3.1-8b 1338.55 1 8
Llama-3.1-70b 3580.30 6 16
Phi-3-medium 4121.94 1 8
Gemma-2-27b 35563.46 1 1
Yi-1.5-34b 3516.06 1 4
Internlm-2.5-7b 1446.18 1 8
Table 23: The total running time of different LLMs, the number of GPUs used, and the batch size.
Models Single batch running time
Qwen-1.5-110b 2763.37
Qwen-2-7b 182.90
Qwen-2-72b 2683.44
Mixtral-8x22b 4836.46
Mixtral-large-2 3075.74
Llama-3-8b 181.25
Llama-3-70b 233.85
Llama-3.1-8b 167.32
Llama-3.1-70b 223.77
Phi-3-medium 515.24
Gemma-2-27b 35563.46
Yi-1.5-34b 879.02
Internlm-2.5-7b 180.77
Table 24: The runtime per batch for different models. This is calculated by directly dividing the total time from Table 23 by the batch size.
Appendix C Case study
In this section, we provide a case study to illustrate the input and output of our experiment. We demonstrate the model’s behavior using example prompts and their corresponding outputs.
C.1 Yue-TruthfulQA
Figure 6: Yue-TruthfulQA Qwen-1.5-110b
Figure 7: Yue-TruthfulQA Gemma-2-27b-it
C.2 Yue-GSM8K
Figure 8: Yue-GSM8K GPT-4o
Figure 9: Yue-GSM8K Gemma-2-27b-it
C.3 Yue-TRANS
Figure 10: Yue-TRANS GPT-4o
Figure 11: Yue-TRANS Qwen-2-72b-Instruct
C.4 Yue-ARC-C
Figure 12: Yue-ARC-C Claude-3.5
Figure 13: Yue-ARC-C ERNIE-Tiny-8k
C.5 Yue-MMLU
Figure 14: Yue-MMLU Qwen-2-72b-Instruct
Figure 15: Yue-MMLU Mixtral-8x22b-Instruct